# ReNF: Rethinking the Design Space of Neural Long-Term Time Series Forecasters

### v4

@@ 2.5 Learning Objective
Our BDO framework uniquely generates multiple outputs of varying lengths, enabling us to apply this loss at each forecasting stage. This hierarchical supervision allows us to fully leverage the label information at multiple scales, encouraging the model to build causally structured and homogeneous representations rather than the disconnected ones typical of standard DO forecasts. In the words of our MNFT, this multi-level supervision provides richer feedback and mutual information from the observed labels, thereby constraining the solution space more effectively.
In the BDO paradigm, the quality of early, short-term forecasts is critical, as errors at these stages may still propagate and degrade the performance of subsequent long-term forecasts. To ensure a stable foundation, we therefore place heavier weights on the losses computed at earlier stages (i.e., for shorter forecast horizons). The complete loss function is expressed as follows:
@@
where $\|\cdot\|_{1}$ denotes the $l_{1}$ norm, $\gamma$ and $\alpha$ is predifined coefficients. $\text{Freq}(\cdot)$ represents the discrete fourier transform and $\hat{Y}_{f}^{(n)}$ denotes the $n$-th sub-forecast corresponding to the $n$-th piece of $X_{f}$. A more in-depth analysis of this loss function is provided in Appendix D.
@@ 2.6 Smoothing the learning for time series
Figure 5: Variation of valid and test loss before and after applying EMA smoothing. The valid loss and test loss are not consistent during the learning process of NFs without smoothing.
Notably, training on standard benchmark datasets is often unstable, partly due to sparse or redundant information. The non-stationary nature of real-world time series, combined with chronological data splitting, leads to both internal (batch-to-batch) and external (train/validation/test) distribution mismatches. This causes the optimization path to be heterogeneous, rendering the learning process for NFs ineffective. Specifically, internal data redundancy can cause the model to overfit repeated or spurious patterns (liu2024timebridge), harming generalization, while external distribution shift creates inconsistencies across the different phases of training. While normalization techniques like RevIN address the train-test shift, a critical problem remains between validation and testing.
On many datasets, particularly those with smaller volumes, the validation and test losses exhibit significantly different, and at times conflicting, dynamics during training. This issue invalidates the early stopping criterion: learning steps become ineffective, and suboptimal models are saved, preventing a true assessment of a model’s capabilities. This is especially problematic when comparing models of varying complexities, which naturally have different optimization trajectories.
To mitigate these unexpected effects, we propose smoothing the training trajectory by employing a ”shadow” model for evaluation. We achieve this efficiently using an Exponential Moving Average (EMA) to track the parameters of the online model, a technique proven effective in self-supervised learning in self-supervised learning (jepa) and generative models (song2020scoreSDE). Specifically, let $\theta$ be the parameters of the online model being trained, and $\theta^{\prime}$ be the parameters of the shadow model. Then the shadow model’s parameters are upgraded as follows at each iteration.
@@
where $\alpha$ is the EMA decay rate. This shadow model is then used for all evaluations. As shown in Fig. 5, this technique effectively smooths the learning curves and mitigates the inconsistency between validation and test performance. By providing a more stable and reliable training signal, EMA prolongs the effective learning period and enables the model to converge to better local minima, achieving improved generalization.
@@ 3 Experiment
Baselines and Datasets. We compare our model with a suite of recent state-of-the-art methods, including TimeBridge (liu2024timebridge), DUET (qiu2025duet), TimeDistill (ni2025timedistill), Timer-XL (timerxl), iTransformer (liu2023itransformer), TimeMixer (wang2024timemixer), PatchTST (patchtst), Crossformer (zhang2023crossformer), and Dlinear (dlinear). We use widely used benchmark datasets in this area, including electricity (ETTh1, ETTh2, ETTm1, ETTm2, Electricity), environment (Weather), energy (Solar-Energy), and transportation (Traffic). Detailed descriptions of these datasets are provided in Appendix B.
Setups. All experiments were conducted on a single NVIDIA 4090 GPU with 24GB of memory, using the Adam optimizer (kingma2014adam) and a fixed random seed of 2021 for reproducibility. Results for all baseline models were reproduced using their official source code and optimal configurations. For our model, ReNF, we searched for the optimal learning rate in the range from 0.0001 to 0.005, the EMA decay rates in the range from 0.99 to 0.999, and the number of layers from 2 to 8. We apply ReNF-$\alpha$ to ETTh1 and ETTh2 datasets, and ReNF-$\beta$ to others. The look-back window of ReNF is searched over $\{336,512,720\}$ for the best performance. The drop_last bug is corrected following the TFB benchmark (qiu2025duet).
@@ 3.1 Main Result
The results, presented in Table 2, demonstrate a powerful conclusion: without resorting to complex, specialized modules for multi-resolution, periodicity, or cross-variate dependencies, a simple MLP-based model can achieve exceptional performance. By focusing instead on fundamentally improving the forecasting paradigm and stabilizing the training process,
@@
our MLP model-ReNF sets a new state-of-the-art. Overall, ReNF surpasses all leading 2024 models by a significant margin and outperforms even the most competitive 2025 SOTA methods in almost all cases. This provides strong evidence for the effectiveness of our proposed techniques.
However, we do not claim that existing specialized techniques are redundant. Rather, as discussed in Sec.2.6, many of these architectural designs were evaluated within unstable training frameworks, suggesting that their true effects may need tedious re-evaluation, which may be a promising future work. For instance, on the complex Traffic dataset, while ReNF shows significant improvement over many baselines, it does not surpass the MSE score of the large Transformer-based model, TimeBridge. We compare the model complexity in FLOPs and params between the most competitive models in the Table 1, showing that ReNF reduces 20x the complexity in FLOPs compared to TimeBridge with the Traffic dataset. This suggests that more complex models might still work better on datasets with high non-linearity.
Models
ReNF TimeBridge DUET TimeDistill Timer-XL iTransformer TimeMixer PatchTST Crossformer DLinear
@@
Table 2: Results of long-term forecasting of hyperparameter searching. All results are averaged across four different prediction lengths: $\{96,192,336,720\}$. The best and second-best results are highlighted in red and blue, respectively. Full results are listed in Appendix E.
@@ 3.2 Ablation Study
First, on smaller or less stable datasets like ETTh2, EMA mitigates spurious overfitting and stabilizes the learning curves. This provides a more reliable signal for early stopping and prolongs the effective training period. Second, on large, high-quality datasets such as Electricity, the choice of EMA decay rate can influence performance. By selecting an appropriate rate, the generalization ability of the NF can be substantially enhanced. Full numerical results for all datasets are available in the Appendix.
@@
Furthermore, it is shown clearly in Fig. 8(a) that the performance of ReNF with the BDO strategy can consistently improve as K increases. In stark contrast, the performance of the DO model stagnates or degrades with added depth. This difference highlights BDO’s ability to effectively utilize deeper architectures. This meaningful phenomenon suggests that our paradigm may unlock a new, more effective scaling law for LTSF models. Diverse examples are shown in the Table 3 and the Appendix E.
3.3 Empirical Evaluation of the Proposition
@@
To empirically evaluate our Multiple Neural Forecasting Proposition, we implement the post-combination strategy described in Sec. 2.2. This requires defining a combination function, $g_{c}$. While finding an optimal solution without future knowledge is difficult, we can establish a theoretical upper bound on performance. Given the ground truth labels, a trivial yet optimal strategy is to select, at each timestep, the forecast value from among all candidates that has the smallest error. By applying this ”oracle” combination, we aim to explore two points: 1) to quantify the potential accuracy improvement achievable through post-combination, thereby establishing a dynamic empirical bound for our forecasting model; and 2) to verify that this empirical bound behaves in a manner consistent with the intuitions of the MNFP 2.2.
@@
Furthermore, Fig. 8(b) shows how this empirical bound varies with the number of sub-forecasts $K$. The bound consistently decays as $K$ increases, even when the model’s performance saturates. This result aligns perfectly with the core insight from MNFP: the theoretical performance limit improves as the number of candidate forecasts grows.
@@
In a nutshell, our evaluation of the MNFP clarifies two primary roles for Neural Networks in this domain. First, NNs should be leveraged to learn a powerful post-combination function over multiple candidate forecasts. Second, more powerful base forecasters must be developed to better approximate the true data-generating distribution, directly reducing the predictive bias $b$ as identified in our proposition 2.2.
@@
We propose a fundamental proposition to support the mechanism of multiple forecasts. Based on this proposition, we introduce a novel paradigm that combines the advantages of current forecasting methods, enabling more accurate predictions in a hierarchical representation space. To remedy training instability, we leverage EMA to smooth the learning process of neural forecasters. Our empirical results show that with these techniques, a simple MLP can outperform recent SOTA models with significantly less complexity. The value of our proposition is also evaluated and confirmed by experimental results, clearly identifying the role of the Neural Network in building the Long-Term Time Series forecasters.
@@ 5 Reproducibility statement
We have put a lot of effort into ensuring the reproducibility of this work, including the code in the anonymous repository , a complete data description in the Appendix B, and clear explanations of the assumptions and proof of the proposition in the Appendix C.
Supplementary Material
Appendix A Related Work
We note that ReNF is not the first work to address LTSF using multiple short-term forecasts. A pioneering and highly relevant study by (nguyen2004multiple) proposed a machine called MNN, which generates long-term forecasts using multiple neural networks, each responsible for a specific interval of the output. The essential distinction between MNN and our approach is that we generate sub-forecasts recursively, with the explicit goal of injecting causality into the DO strategy. In contrast, each sub-network in MNN functions as a standalone AR(1) model. This difference stems from our distinct motivation: whereas MNN was designed primarily to mitigate the error accumulation of one-step-ahead NN models, our work begins with a foundational forecasting proposition. Based on this proposition, we leverage modern deep learning techniques to effectively utilize multiple sub-forecasts, achieving strong empirical results on a wide range of real-world datasets, which constitutes one of our main contributions.
Our work also connects to classic ensemble methods in machine learning (mohri2018foundationsofml). For instance, the recursive workflow of ReNF is analogous to gradient-boosted regression models (chen2016xgboost), which construct a strong predictor from multiple weak ones. Additionally, the collaborative training can be viewed as a form of bagging (breiman1996bagging), which effectively resamples the data to train diverse forecasters. In this context, an even more recent study (emsenble4trans) re-certified the benefits of such ensemble methods for enhancing Transformer-based NFs, providing further empirical support for the direction of our research.
Forecast Combinations. Forecast combination is a classic technique for improving forecast accuracy and robustness by leveraging the diverse strengths of multiple models (combiningreview). In this work, we extend this concept (bates1969combination) to the domain of deep learning for long-term time series forecasting. Rather than combining distinct, parallel forecasters into a hybrid model (hybrid), our framework achieves this goal efficiently within a single, structured approach for generating and implicitly combining forecasts within a single neural network. By recursively stacking sub-forecasts, similar in spirit to N-BEATS (nbeats), we ensure that the representation learning is both deep and diverse, allowing the LSTF task to benefit from the full power of modern machine learning.
Appendix B Details of Datasets
Dataset Dim Prediction Length Split Frequency domain
@@ Appendix B Details of Datasets
Table 5: Descriptions of multivariate time series datasets used in this research. The dim column represents the number of variates, and the split column specifies the train-validate-test splitting ratio for each dataset.
The ETTh1, ETTh2, ETTm1, and ETTm2 datasets record the temperature of electricity transformers every hour and every 15 minutes. The Weather dataset contains 21 weather information, measured every 10 minutes in Germany. The Electricity dataset records the amount of electricity used by 321 customers every hour. The ILI dataset records how many patients there are and how many have influenza-like illnesses (ILI) every week. The Solar dataset records how much electricity is produced by solar power stations every 10 minutes, from 137 solar power stations, in 2006. The Traffic dataset records how busy the roads are in San Francisco, every hour, from 862 sensors on the freeway. We also include the PEMS04 dataset, which contains information about the traffic network in California, collected every 5 minutes. This is often used to make predictions about the movement of traffic.
Appendix C Proof of the approximate proposition
We restate and emphasize the preconditions of our proof:
  • (A1) The elements in the time series are bounded by a positive real constant $\lambda$. (We also assume the observed series is normal and noiseless for simplicity.)
  • (A2) The NFM generates $c$ candidate series $\{\hat{Y}^{(i)}_{f}\}_{i=1}^{c>1}$ under various $\theta^{(i)}$ or random states $\gamma^{(i)}$.
  • (A3) The predictive bias of the NFM is bounded: $\sup_{t}{|{\hat{\mu}}_{t}-\mu_{t}|}=b$

### v5

@@ 2.4 Learning Objective
Our BDO framework uniquely generates multiple outputs of varying lengths, enabling us to apply this loss at each forecasting stage. This hierarchical supervision allows us to fully leverage the label information at multiple scales, encouraging the model to build causally structured and homogeneous representations rather than the disconnected ones typical of standard DO forecasts.
To ensure a stable foundation as well as the diversity of different forecasts, we place heavier weights on the losses of earlier short-term sub-forecasts. The complete loss function is expressed as follows:
@@
where $\|\cdot\|_{1}$ denotes the $l_{1}$ norm, $\gamma$ and $\alpha$ is predifined coefficients. $\text{Freq}(\cdot)$ represents the discrete fourier transform and $\hat{Y}_{f}^{(n)}$ denotes the $n$-th sub-forecast corresponding to the $n$-th block of the model. A more in-depth analysis of this loss function is provided in Appendix C.
@@ 2.5 Smoothing the learning for time series
Figure 5: Variation of valid and test loss with and without applying EMA weight averaging. The valid loss and test loss are not consistent during the learning process of NFs without smoothing.
On many time series datasets, particularly those with smaller volumes, the validation and test losses exhibit significantly different and, at times, conflicting dynamics during training. This issue invalidates the early stopping criterion: learning steps become ineffective, and suboptimal models are saved, preventing a true assessment of a model’s capabilities. This is especially problematic when comparing models of varying complexities, which naturally have different optimization trajectories.
To mitigate these unexpected effects, we propose smoothing the training trajectory by employing an averaged model for evaluation. We achieve this efficiently using an Exponential Moving Average (EMA) to track the parameters of the online model (swa). While this has been a standard technique in self-supervised learning (moco) and generative modeling (song2020scoreSDE), its critical role in mitigating the specific distribution shifts found in LTSF has been largely underexplored. We empirically establish it as a necessary protocol for robust performance. Specifically, let $\theta$ be the parameters of the online model being trained, and $\theta^{\prime}$ be the parameters of the shadow model. Then the shadow model’s parameters are upgraded as follows at each iteration.
@@
where $\alpha$ is the EMA decay rate. This shadow model is then used for all evaluations. As shown in Fig. 5, this technique effectively smooths the learning curves and mitigates the inconsistency between validation and test performance. By providing a more stable and reliable training signal, EMA prolongs the effective learning period and allows the model to converge to better local minima, achieving improved generalization.
@@ 3 Experiment
Baselines and Datasets. Our empirical evaluation is designed to first establish the core strengths of our method in its intended domain and then to verify its generality across a wider range of forecasting tasks. Accordingly, the main text concentrates on established multivariate long-term forecasting benchmarks (ETTh, Electricity, Weather, etc.), where the hierarchical, variance-mitigating structure of BDO is particularly effective. To validate the versatility of our approach, we then present a comprehensive suite of supplementary experiments in the Appendix. These evaluations test our proposals on distinct domains, including diverse short-term forecasting from Kaggle (yue2025olinear), spatio-temporal traffic data (PEMS), and classical univariate series (M4 competition (wang2024timemixer)). A complete description of these datasets can be found in Appendix A.
For the multivariate long-term forecasting, we compare our model with a suite of recent state-of-the-art methods, including TimeBridge (liu2024timebridge), DUET (qiu2025duet), TimeDistill (ni2025timedistill), Timer-XL (timerxl), iTransformer (liu2023itransformer), TimeMixer (wang2024timemixer), PatchTST (patchtst), Crossformer (zhang2023crossformer), and Dlinear (dlinear).
Setups. All experiments were conducted on a single NVIDIA 4090 GPU with 24GB of memory, using the Adam optimizer (kingma2014adam) and a fixed random seed of 2021 for reproducibility. Results for all baseline models were reproduced using their official source code and optimal configurations. For our model, ReNF, we searched for the optimal learning rate in the range from 0.0001 to 0.005, the EMA decay rates in the range from 0.99 to 0.999, and the number of layers from 2 to 8. We apply ReNF-$\alpha$ to ETTh1 and ETTh2 datasets, and ReNF-$\beta$ to others. The look-back window of all models is searched over $\{336,512,720\}$ for the best long-term forecasting performance following the standard benchmark configuration (qiu2025duet).
@@ 3.1 Main Result
The results, presented in Table 2, demonstrate a critical finding: a well-executed, straightforward MLP-based architecture can still outperform complex state-of-the-art models. Notably, we achieve these gains without relying on intricate mechanisms such as multi-resolution and periodicity. By focusing instead on fundamentally improving the forecasting principles,
@@
our MLP model-ReNF sets a new state-of-the-art. Overall, ReNF surpasses all leading 2024 models by a significant margin and outperforms even the very competitive 2025 SOTA methods in almost all cases. This manifestation of contrast provides strong evidence for the effectiveness of our proposed techniques.
However, we do not claim that existing specialized techniques are redundant. Rather, as discussed in Sec.2.5, many of these architectural designs were evaluated within unstable training frameworks, suggesting that their true effects may need tedious re-evaluation, which may be a promising future work. For instance, on the Traffic dataset, while ReNF shows significant improvement over many baselines, it does not surpass the MSE score of the large Transformer-based model, TimeBridge. We compare the model complexity between the most competitive models in the Table 1, showing that ReNF reduces 20x the complexity in FLOPs compared to TimeBridge with the Traffic dataset. This suggests that more complex models might still work better on datasets with high volume.
Models ReNF TimeBridge DUET TimeDistill Timer-XL iTransformer TimeMixer PatchTST Crossformer DLinear
@@
Table 2: Results of long-term forecasting of hyperparameter searching. All results are averaged across four different prediction lengths: $\{96,192,336,720\}$. The best and second-best results are highlighted in red and blue, respectively. Full results are listed in Appendix D.
@@ 3.2 Ablation Study
First, on smaller or less stable datasets like ETTh2, EMA mitigates spurious overfitting and stabilizes the learning curves. This provides a more reliable signal for early stopping and prolongs the effective training period. Second, on large, high-quality datasets such as Electricity, the choice of EMA decay rate can influence performance. By selecting an appropriate rate, the generalization ability of the NF can be substantially enhanced. Full numerical results for all datasets are available in the Appendix E.3. These results show clearly the advantages of leveraging the model weight averaging technique and strongly support our claim in addressing the issue of valid-test inconsistency when training a neural forecaster for time series.
@@
Furthermore, it is shown clearly in Fig. 8(a) that the performance of ReNF with the BDO strategy can consistently improve as $K$ increases. In stark contrast, the performance of the DO model stagnates or degrades with added depth. This difference highlights BDO’s ability to effectively utilize deeper architectures. This meaningful phenomenon suggests that our paradigm may unlock a new, more effective scaling potential for NFs. Diverse examples are shown in the Table 3 and the Appendix D.
3.3 Empirical Evaluation of the Hypothesis
@@
To empirically evaluate our variance reduction hypothesis, we implement the post-combination strategy described in Sec. 2.1. This requires a combination functional, $g_{c}$. While finding an optimal solution in practice is difficult, we can establish a theoretical upper bound on performance for the use of evaluation. Given the ground truth labels, a trivial yet optimal strategy is to select, at each timestep, the forecast value from among all candidates that has the smallest error. By applying this ”oracle” combination, we aim to explore two points: 1) to quantify the potential accuracy improvement achievable through post-combination, thereby establishing a dynamic empirical bound for our forecasting model; and 2) to verify that this empirical bound behaves in a manner consistent with the intuitions of the VRH 2.1.
@@
Furthermore, Fig. 8(b) shows how this empirical bound varies with the number of sub-forecasts $K$. The bound consistently decays as $K$ increases, even when the model’s performance saturates. This result aligns perfectly with the core insight from VRH: the theoretical performance limit improves as the number of candidate forecasts grows.
@@
In a nutshell, our evaluation of the VRH clarifies two primary roles for Neural Networks in this domain. First, NNs should be leveraged to learn a powerful post-combination function over multiple candidate forecasts. Second, more powerful base forecasters must be developed to better approximate the true data-generating distribution, directly reducing the predictive bias $b$ as identified in our proposition 2.1.
4 Related Work
Ensemble. Our work connects to classic ensemble methods in machine learning (mohri2018foundationsofml). For instance, the recursive workflow of ReNF is analogous to gradient-boosted regression models (chen2016xgboost), which construct a strong predictor from multiple weak ones. Additionally, the collaborative training can be viewed as a form of bagging (breiman1996bagging), which effectively resamples the data to train diverse forecasters. In this context, an even more recent study (emsenble4trans) re-certified the benefits of such ensemble methods for enhancing Transformer-based NFs, providing further empirical support for the direction of our research.
Forecast combination is a classic ensemble technique for improving forecast accuracy and robustness by leveraging the diverse strengths of multiple models (combiningreview). In this work, we revisit this concept (bates1969combination) in the domain of deep learning for long-term time series forecasting. Rather than combining distinct, parallel forecasters into a hybrid model (hybrid), our framework achieves this goal efficiently within a single, structured approach for generating and implicitly combining forecasts within a single neural network.
Our approach shares a conceptual lineage with deep stacking architectures (stacking) like N-BEATS (nbeats) and N-HiTS (nhits), which also synthesize a final forecast from multiple sub-modules. However, the fundamental motivation and aggregation mechanism differ. While N-BEATS and N-HiTS rely on additive residual decomposition—where subsequent blocks focus on correcting the errors (or specific frequency components) of previous blocks—our BDO paradigm employs recursive input concatenation. Instead of summing independent components, ReNF generates forecasts in a curriculum-style progression (from short-term to long-term horizons). By explicitly feeding earlier sub-forecasts back into the input space of subsequent layers, ReNF enforces a sequential dependency and allows the model to iteratively refine its understanding of the temporal trajectory, rather than performing a static multi-scale decomposition.
Short to Long Forecasting. We note that ReNF is not the first work to address LTSF using multiple short-term forecasts. A pioneering and highly relevant study by (nguyen2004multiple) proposed a machine called MNN, which generates long-term forecasts using multiple neural networks, each responsible for a specific interval of the output. The essential distinction between MNN and our approach is that we generate sub-forecasts recursively, with the explicit goal of injecting causality into the DO strategy. In contrast, each sub-network in MNN functions as a standalone AR(1) model. This difference stems from our distinct motivation: whereas MNN was designed primarily to mitigate the error accumulation of one-step-ahead NN models, our work begins with a foundational forecasting proposition. Based on this proposition, we leverage modern deep learning techniques to effectively utilize multiple sub-forecasts, achieving strong empirical results on a wide range of real-world datasets, which constitutes one of our main contributions.
@@ 4 Related Work
We locate some fundamental issues in the current design of neural forecasters. By proposing a Variance Reduction Hypothesis, we introduce a novel forecasting paradigm that combines the advantages of current forecasting methods and the principle of forecast combination, enabling more accurate predictions in a hierarchical structured representation space. We identify the weight averaging technique (e.g., EMA) as a critical, yet often overlooked, component for resolving the validation-test discrepancy in neural forecasting, serving to smooth the volatile learning process inherent to time series data. Our empirical results show that with these techniques, a simple MLP can still outperform recent SOTA models. The value of our hypothesis is also evaluated and confirmed by experimental results, clearly identifying the role of the Neural Network in building the Long-Term Time Series forecasters.
@@ 6 Reproducibility statement
We have put a lot of effort into ensuring the reproducibility of this work, including the code in the anonymous repository https://anonymous.4open.science/r/ReNF-A151, a complete data description in the Appendix A, and clear explanations of the assumptions and proof of the proposition in the Appendix B.1.
Appendix
List of Appendices
Appendix A Details of Datasets
Dataset Variate Prediction Length Split Frequency domain
@@ Appendix A Details of Datasets
NASDAQ 12 24, 36, 48, 60 (7, 1, 2) Daily Finance
SP500 12 24, 36, 48, 60 (7, 1, 2) Daily Finance
Carsales 12 24, 36, 48, 60 (7, 1, 2) Daily Market
Website 12 24, 36, 48, 60 (7, 1, 2) Daily Web
Power 12 24, 36, 48, 60 (7, 1, 2) Daily Energy
METR-LA 207 24, 36, 48, 60 (7, 1, 2) 5min Transportation
Table 5: Descriptions of multivariate time series datasets used in the table 6 and table 7. The Variate column represents the number of variates, and the split column specifies the train-validate-test splitting ratio for each dataset.
The ETTh1, ETTh2, ETTm1, and ETTm2 datasets record the temperature of electricity transformers every hour and every 15 minutes. The Weather dataset contains 21 pieces of weather information, measured every 10 minutes in Germany. The Electricity dataset records the amount of electricity used by 321 customers every hour. The Solar dataset records how much electricity is produced by solar power stations every 10 minutes, from 137 solar power stations, in 2006. The Traffic dataset records how busy the roads are in San Francisco, every hour, from 862 sensors on the freeway.
In addition, we also evaluate the short-term time series forecasting on eight multivariate datasets that are used in the study (yue2025olinear). The METR-LA database was populated with traffic network data for Los Angeles during the spring of 2012, specifically from March to June. This data was collected at an interval of five minutes. The NASDAQ includes the daily NASDAQ index and key economic indicators from 2010 to 2024. The SP500 records daily SP500 index data (e.g., opening price, closing price, and trading volume) from January 1993 to February 2025. CarSales collects daily sales of 10 vehicle brands (e.g., Toyota, Honda) in the U.S. from January 2005 to June 2023. The data are compiled from the Vehicle Sales dataset on Kaggle. Power contains daily wind and solar energy production (in MW) records for the French grid from April 2020 to June 2023. The data are compiled from the Wind and Solar Daily Power Production dataset on Kaggle. The website contains six years of daily visit data (e.g, first-time and returning visits) to an academic website, spanning from September 2014 to August 2020.
Appendix B Theoritical Analysis
B.1 Proof of the Hypothesis